OpenStudy (anonymous):

A restaurant serves 8 fish entrees, 12 beef and 10 poultry. If customers select from these entrees randomly, what is the expected number of fish entrees ordered by the next four customers?

OpenStudy (anonymous):

There are 8 fish entrees--so the expected number of fish entrees ordered by the next 4 customers is 960?

OpenStudy (anonymous):

...doesn't make sense

OpenStudy (campbell_st):

ok.. a little probability so P(fish) = 8/30 P(fish, fish, fish , fish) = 8/30 x 8/30 x 8/30 x 8/30

OpenStudy (anonymous):

Okay, that makes more sense. Thank you